The low pressure side of the 3,300v/220v transformer was opened and the high pressure side was applied with 3,300v. In this case, current 1A flows and the power factor is 0.12.

What would the current flow be if the high pressure side of the transformer was opened and 220v applied to the low pressure side? Also, what is the effective and reactive power?
Answer #1

Real power when LV side is opened

P=VI*pf

P=3300*1*0.12

P=396W

Reactive power

Q=3300*1*0.992

Q=3276.1 Var

Now HV side opened

S=VI

220*I=3300*1

I=15A

P=396W

Q=3276.1 Var